what are the factors of 6r^2+3rs-18s^2
Answer by mananth(16946)

what are the factors of 6r^2+3rs-18s^2
divide by 3
2r^2+rs-6s^2
2r^2+4rs-3rs+6r^2=0
2r(r+4s)-3s(r+4s)=0
(2r-3s)(r+4s)